OpenStudy (anonymous):

Help with this one. (3sqrt3 + sqrt 2) (sqrt3 - sqrt2) expressed as a + bsqrtc

OpenStudy (anonymous):

\[(3\sqrt{3} + \sqrt{2}) (\sqrt{3} - \sqrt{2}) \]

OpenStudy (anonymous):

Expressed in \[a + b \sqrt{c}\] form!

OpenStudy (anonymous):

9 - 3sqrt3*sqrt2 +sqrt3*sqrt2 - 2 = 7 - 2sqrt3*sqrt2

OpenStudy (anonymous):

Foil it out: \[3(\sqrt3)^2-3\sqrt2\sqrt3+\sqrt2\sqrt3-(\sqrt2)^2=9-3\sqrt6+\sqrt6-2=7-2\sqrt6\]

OpenStudy (anonymous):

= 7 - 2sqrt6